A wheelbarrow manufacturer uses handles made of hardwood. Find the amount of wood that is necessary to make 182 handles. The longest dimension shown is the handle length. The following information was provided.
6.0 cubic Ft WheelBarrow
* steel construction with hardwood handles
* 14" tubeless pneumatic tire
* fully rolled edge for added tray strength
* overall dimensions: 61 1/2"L x 27"W x 24.9"H

Answer by stanbon(48552)

You have to assume the handles are 61 1/2 inches long.
182 of them would be 182*(123/2)" = 91*123 =11193 inches of wood
or 932.75 ft of wood
--------
Cheers,
Stan H.